OpenStudy (anonymous):

Please Explain how to do this. A train starts from rest and accelerates at a rate of 1.6m/s^2 for 10s. It runs at constant speed for 70s and slows down for a rate of 3.5m/s^2 until it stops. Find total distance covered.

OpenStudy (anonymous):

Calculate velocity after 10s v = u + at v = 1.6*10 = 16m/s Distance traveled in first 10 sec s = ut + 0.5 at^2 s1 = 0.5*1.6*100 = 80m Distance traveled with constant speed s2 = 16 * 70 = 1120m Distance traveled during braking v^2 = u^2 + 2 as 0 = 256 - 2*3.5*s3 s3 = 36.6 m Total distance is s1 +s2 +s3

OpenStudy (anonymous):

Thanks shadowfax that looks right.